Obtain an expression for mean free path of the molecules in a gas in the first order
approximation. Also, show that mean free path of the molecules in a gas decreases
when all molecules are moving
1
Expert's answer
2020-02-26T10:18:56-0500

It is the average distance traveled by molecule between the successive collisions, which modify it's direction or energy or other particle property.

Let the duration between the collision is t and the distance covered by the molecule is "\bar{v}t" and the diameter of the cylinder is 2d.

so, volume of the cylinder "=\pi d^2 vt"

Average number of collisions z="n_v\pi d^2 vt"

where "n_v" is the number of collisions.

So, mean free path(l) ="\dfrac{vt}{n_v\pi d^2 vt}=\dfrac{1}{n_v\pi d^2}"

we know that "n_v=\dfrac{p}{K_bT}"


So, "l=" "\dfrac{1}{n_v\pi d^2}=\dfrac{K_bT}{p\pi d^2}"

we know that, "v\rightarrow\sqrt{2}v"

So, "z=n_v\pi d^2 vt=n_v\pi d^2 \sqrt{2}vt"

Hence, mean free path (l)"=\dfrac{vt}{n_v\pi d^2 \sqrt{2}vt}=\dfrac{1}{\sqrt{2}n_v\pi d^2}"

"l=" "\dfrac{K_b T}{\sqrt{2}p\pi d^2}"

As the number of molecules are getting increase, the molecules will become closer to each other, therefore they more likely to run each other so mean free path decreases.